Assume the cost of a company picnic is described by the function
P(n)=(1/2)n^2-10n +80
where n represents the number of employees and family members attending the picnic and P (in dollars) represents the cost of the picnic. How many employees and guests in attendance produce a Minimum Cost? What is the Minimum cost for this event?

Answer by stanbon(75887)

P(n)=(1/2)n^2-10n +80
You have a quadratic with a=1/2, b=-10
Minimum is at n=-b/2a = 10/(1)= 10 people attending.
P(10)= (1/2)100-10(10)+80
P(10)= 50-20= $30 (the cost of the party)
Cheers,
Stan H.